A Business Firms sells a good at the price of rs 450. The firm has decided to reduce the price of good to rs 350. Consequently , the quantity demanded for the good rose from 25,000 units to 35,000 units. Calculate the price elasticity of demand
1
Expert's answer
2020-11-27T14:17:19-0500

"Arc elasticity =\frac{ Percentage change in demand }{ Percentage change in price}"


"Percentage change in demand = \frac{35000-25000}{(35000+25000)/2} \times100"


"= \frac{10000}{30000}\times100"

="33\frac{1}{3}\%"


"Percentage change in price = \frac{350-450}{(350+450)/2} \times100"


"Percentage change in price =- \frac{100}{400} \times100"


"= -25\%"


Therefore Arc elasticity "= \frac{percentage change in demand}{percentage change in price}"


"=\frac{33.33} {-25}"


"= - 1\frac{1}{3}"